# mode=buildnew

### v2

@@ Abstract
We address the online combinatorial choice of weight multipliers for multi-objective optimization of many loss terms parameterized by neural works via a probabilistic graphical model (PGM) for the joint model parameter and multiplier evolution process, with a hypervolume based likelihood promoting multi-objective descent. The corresponding parameter and multiplier estimation as a sequential decision process is then cast into an optimal control problem, where the multi-objective descent goal is dispatched hierarchically into a series of constraint optimization sub-problems. The subproblem constraint automatically adapts itself according to Pareto dominance and serves as the setpoint for the low level multiplier controller to schedule loss landscapes via output feedback of each loss term. Our method is multiplier-free and operates at the timescale of epochs, thus saves tremendous computational resources compared to full training cycle multiplier tuning. It also circumvents the excessive memory requirements and heavy computational burden of existing multi-objective deep learning methods. We applied it to domain invariant variational auto-encoding with 6 loss terms on the PACS domain generalization task, and observed robust performance across a range of controller hyperparameters, as well as different multiplier initial conditions, outperforming other multiplier scheduling methods. We offered modular implementation of our method, admitting extension to custom definition of many loss terms.
@@ 1 Introduction
In many deep learning fields including domain generalization (Gulrajani and Lopez-Paz, 2020), the loss function for neural networks amounts to a summation of multiple terms, with a multiplier weighting each term. As the number of loss terms increases, the space of choice for those multipliers grows dramatically. In addition, when the validation set is drawn from a different distribution than the target domain, the validation performance does not faithfully reflect the target domain performance (Chen et al., 2022b) and can easily reach and maintain a saturated value during training. All these factors complicate and add difficulties to the process of model selection using conventional hyperparameter tuning (Probst et al., 2019) methods.
Instead of optimizing a fixed weighted summation of many neural network loss terms using a single objective optimization algorithm which relies on heavy hyperparameter tuning of the weight multipliers, multi-objective optimization on these loss terms has raised attention in the deep learning community recently (Sener and Koltun, 2018; Lin et al., 2019; Mahapatra and Rajan, 2021; Chen et al., 2022b). Previous works in this direction (Mahapatra and Rajan, 2021; Chen et al., 2022b) addresses the problem via careful choice of multipliers at each optimization iteration instead of fixed for the whole training cycle, but suffers from excessive memory requirement and heavy extra computation burden, which hinders their applicability to large dataset and large scale neural networks Chen et al. (2022b). Furthermore, it is restricted to Exact Pareto optimization (Mahapatra and Rajan, 2021) where the solution should be aligned to a user input preference vector, while how to set that is not clear in the general case. Their performance further relies on careful tuning of learning rate of single objective low level optimization algorithms Chen et al. (2022b). Single objective optimization algorithms like ADAM (Kingma and Ba, 2015) were analyzed predominantly with convex optimization theory (Kingma and Ba, 2015; Reddi et al., 2019), which does not hold globally (Milne, 2019). From a control theory point of view, these low level optimization algorithms drive the model parameters iteratively, which forms a dynamic system. Due to lack of comprehensive theoretical underpinnings of the loss function property with respect to model parameters, this dynamic system has unreliable open-loop behavior.
@@
Inspired by feedback control theory (Doyle et al., 2013) and feedback optimization in power systems (Picallo et al., 2020; He et al., 2023; Hauswirth et al., 2024), we treat the dynamic system of model parameters driven by the low level optimization algorithm as an uncontrolled plant, with model parameter as state and the many loss terms as output. We use state dependent multiplier induced loss landscapes as control input and designed a hierarchical control structure (Dietterich et al., 1998; Kulkarni et al., 2016; Sun et al., 2020) to drive the system towards a multi-objective descent goal, without modifying the internal dynamic of low level neural network optimization algorithms.
This hierarchical control structure allows for a more refined and flexible approach to high-level optimization of neural networks, by automatically balancing the trade-off between a task-specific empirical loss and various regularization losses, which we detail in  Equation 1. We depicted our idea in the control diagram in Figure 1 where the motivation of choosing multiplier induced landscape scheduling as control input is illustrated in Figure 4.
@@
We demonstrated that our method removes the combination curse of dimensionality of multi-dimensional multipliers, via training domain invariant variational auto-encoding (Ilse et al., 2020) on the PACS dataset. Our method only has a small number of hyperparameters for our controllers while achieving rubust performance when changing them. The baselines include different warmup multiplier scheduling strategies, which still need explicit specification of multiplier ultimate values. Their performance varies drastically with changes in those ultimate values. We attribute the improved performance of our method to the automatic trade-off of different loss terms. Our major contributions are:
  • We proposed a probabilistic graphical model (PGM) for depicting the joint multiplier and model parameter adaptation process with a likelihood function imposed, promoting multi-objective descent of each loss term, which can be used to described general multi-objective deep learning. We proposed to transform the problem of parameter inference of this PGM into a sequential decision process through an optimal control formulation.
  • We proposed to solve the sequential decision process for the multi-objective descent goal via breaking the goal into a series of constraint optimization sub-goals which forms a hierarchical controller that automatically adjusts the value of both the multiplier and the model parameters during the training process. The constraint bound updates itself via Pareto dominance. The multi-objective descent of the empirical risk and multidimensional penalty (regularization) loss in Equation 1 offers additional model selection criteria along with validation set performance. Compared to Bayesian optimization, our controller operates at the timescale of model parameter dynamic epoch instead of a complete training, which saves tremendous computational resources. Compared to existing multi-objective deep learning methods, it circumvent excessive memory requirements and heavy extra computational burden.
@@
  • When treating the neural network model parameter dynamic system driven by low level optimization algorithm as uncontrolled plant, we proposed to use state dependent multiplier induced loss landscapes as control input (see Figure 4) and offered initial theoretical analysis of the multi-objective descent behavior of our closed loop system . We offered modular implementation of our method  [^1] https://github.com/marrlab/DomainLab/tree/mhof and showed robustness of our method with respect to different multiplier initial conditions and controller hyperparameters for domain generalization on the PACS dataset with domain invariant auto-encoding.
@@ 2.1 Structural Risk Minimization
In Equation 1, we consider loss $L$ with training data $\mathcal{D}_{tr}$ and model parameters $\theta$. $\ell(\theta,\mathcal{D}_{tr})$ represents the empirical risk and $R(\theta,\mathcal{D}_{tr})$ represents the regularization term of the loss function, also referred to as the penalty function and thus $L$ the penalized loss. $\mathbb{R}_{+}^{d}$ is the $d$ dimensional positive octant. We refer to the multiplier $\mu$ as the penalty multiplier, as a special case of general hyperparameters (e.g. learning rate) in deep learning.
Notation
In the following sections, for brevity, we omit the implicit dependence of $\ell(\cdot)$ and $R(\cdot)$ on $\mathcal{D}_{tr}$ and $\theta$. We also write $R(\theta,\cdot)$ where we need an explicit dependence on $\theta$ but use $\cdot$ only to represent data $\mathcal{D}_{tr}$ and/or other arguments. We use subscript of $R(\cdot)$ to indicate the component of $R(\cdot)$ corresponding to that subscript, e.g., if $\mu=[\beta,\gamma]$, then $R_{\beta}(\cdot)$ corresponds to the component of $R(\cdot)$ associated with $\beta$. We use $X,Y\sim\mathcal{D}$ to represent a random variable $X$ and its supervision signal $Y$ from distribution $\mathcal{D}$ and we use the same symbol $\mathcal{D}$ to denote a corresponding dataset as well. Let $x$ be an instance of $X$, so is $y$ for $Y$.
We use superscript $k$ in bracket to index the optimization iteration, which by default corresponds to one epoch for our method (it can also correspond to one training cycle in Equation 2 to iteratively achieve Equation 3 for conventional hyperparameter optimization, see Remark 3.1).
Definition 2.1.
When optimizing Equation 1 with multiplier $\mu$ iteratively (e.g. with gradient based methods like ADAM), the model parameter evolution dynamic $\theta^{(k+1)}=f_{\theta}(\mu,\theta^{(k)},\mathcal{D}_{tr})$ brings $\theta^{(k)}$ to its next value $\theta^{(k+1)}$. $\theta^{(k+1)}$ then becomes the initial value for further iterations. For notational simplicity, we replace the explicit dependence of this dynamic on training data $\mathcal{D}_{tr}$, multiplier $\mu$ (potentially varying value for each iteration) and/or even the initial model parameter $\theta$ with dots. This convention results in notations $f_{\theta}(\theta,\cdot)$, $f_{\theta}(\mu,\cdot)$ and $f_{\theta}(\cdot)$. We also write $\theta^{+}=f_{\theta}(\mu,\theta,\cdot)$ and $\theta^{+}=f_{\theta}(\theta,\cdot)$ (potentially varying $\mu$ for each iteration) to represent the model parameter dynamic system without explicitly using $k$.
@@ 2.2 Multi-objective optimization
Definition 2.4.
If $\theta_{1}\sim\equiv\theta_{2}$ where $\theta_{1}\in s\mathcal{R}(\theta^{(0)},f_{\theta}(\cdot))$ and $\theta_{2}\in s\mathcal{R}(\theta^{(0)},f_{\theta}(\cdot))$ as in Definition 2.3, we define $\theta_{2}\in\mathcal{C}_{\ell(\cdot),R(\cdot)}\left(\theta_{1},f_{\theta}(\cdot),\theta^{(0)}\right)$ the map from a representing element $\theta_{1}$ to its equivalence class, which contains $\theta_{2}$. We use $\mathcal{E}_{\ell(\cdot),R(\cdot)}\left(\theta_{1},f_{\theta}(\cdot),\theta^{(0)}\right)$ to represent the set of $\ell(\cdot),R(\cdot)$ values for each element of $\mathcal{C}_{\ell(\cdot),R(\cdot)}\left(\theta_{1},f_{\theta}(\cdot),\theta^{(0)}\right)$
2.3 Hyperparameter tuning
The multiplier $\mu$ in Equation 1 is a hyperparameter that affects the evolution of the model parameters $\theta$ in the optimization process for $L$ defined in Equation 1, which leads to optimized model parameter $\theta_{\mu,\mathcal{D}_{tr}}$ below, where we omit the potential dependence of $\theta_{\mu,\mathcal{D}_{tr}}$ on its initial condition $\theta^{(0)}$ for notational simplicity.
@@ 2.3 Hyperparameter tuning
Hyperparameter optimization aims to adjust the multiplier $\mu$ in alignment with specific performance metrics ${O}$ (e.g. validation set accuracy for a classification task) in Equation 3 evaluated on the validation set $\mathcal{D}_{val}$, which lead to the selected hyperparameter $\mu_{\mathcal{D}_{tr},\mathcal{D}_{val}}$.
@@
To solve Equation 3 iteratively, each iteration relies on a complete training cycle to optimize Equation 2. The selection of $\mu$ depends on validation set $\mathcal{D}_{val}$, which can only be drawn from a different distribution compared to the target domain for the domain generalization problem (Section 2.4). Additionally, the performance metric $O$ on the in-domain validation set can reach and maintain a saturated value during training with neural networks of high expressive power.
2.4 Domain generalization
$L(\theta,\mu,x,y,d)=\ell(\cdot)+\mu^{T}R(\cdot)=$ (4)
$\gamma_{y}\mathbb{E}_{q_{\phi_{y}}(z_{y}|x)}[\log q_{\omega_{y}}(y|z_{y})]+$ (5)
$\mu_{recon}\mathbb{E}_{q_{\phi_{d}}(z_{d}|x),q(z_{x}|x),q(z_{y}|x)}\log p_{\theta_{r}}(x|z_{d},z_{x},z_{y})+$ (6)
$[-\beta_{x}KL(q_{\phi_{x}}(z_{x}|x)||p_{\theta_{x}}(z_{x}))]+$ (7)
$[-\beta_{y}KL(q_{\phi_{y}}(z_{y}|x)||p_{\theta_{y}}(z_{y}|y))]+$ (8)
$[-\beta_{d}KL(q_{\phi_{d}}(z_{d}|x)||p_{\theta_{d}}(z_{d}|d))]+$ (9)
$\gamma_{d}\mathbb{E}_{q_{\phi_{d}}(z_{d}|x)}[\log q_{\omega_{d}}(d|z_{d})]$ (10)
In (Ilse et al., 2020), $\mu_{recon}=1.0$ in Equation 6. Additionally, $\gamma_{y}$ and $\gamma_{d}$ are maintained as constants. $\gamma_{y}$ is corresponding to $\ell(\cdot)=\gamma_{y}\mathbb{E}_{q_{\phi_{y}}(z_{y}|x)}[\log q_{\omega_{y}}(y|z_{y})])$ in Equation 5, and $\gamma_{d}$ in Equation 10 is associated with one component of $R(\cdot)$. However, the combined choice of $\gamma_{y}$ and $\gamma_{d}$ significantly influences the generalization performance, as shown by our experimental findings in Section 5.
@@ 3 Methods
In this section, we elaborate on the hierarchical output feedback control strategy for the multidimensional multiplier adaptation in Equation 1 (see Figure 1) towards multi-objective optimization. We propose a Probabilistic Graphical Model (PGM) for the joint model parameter and multiplier adaptation process in Section 3.1, which leads to an optimal control formulation of parameter estimation, extending the classical hyperparameter tuning-based approach to parameter-multiplier co-evolution in Section 3.2.
@@
Figure 1: A control diagram illustrating the hierarchical output feedback optimization process we proposed with multi-objective setpoint adaptation. The uncontrolled plant corresponds to an open loop dynamic $\theta^{+}=f_{\theta}(\mu,\theta,\cdot)$ via optimizing Equation 1 with a low level optimization algorithm (e.g. ADAM) lacking feedback. The $\mu$ controller adjusts the multiplier $\mu$, thus schedules loss landscape (see Illustration in Figure 4), based on the difference $e$ between the setpoint $b$ (defined in Equation 23) and the measured output components $R(\cdot)$, guiding the optimization of model parameters $\theta$ through a feedback loop. The setpoint $b$ is adjusted via feedback from $\ell(\cdot)$ and $R(\cdot)$ (Equation 24a) which forms a higher hierarchy.
3.1 Probabilistic graphical model for parameter and multiplier adaptive optimization process
Figure 2: A probabilistic graphical model for the sequential decision process of joint multiplier and model parameter adaptation in multiple-domain (data-site specific distribution) supervised learning.
@@ 3.1 Probabilistic graphical model for parameter and multiplier adaptive optimization process
Gradient based optimization algorithms for neural network defines a dynamic system for model parameters where the weight multiplier in Equation 1 configures such a dynamic system. Instead of using fixed multipliers (or fixed multipliers’ ultimate values), we propose to choose the multiplier values also dynamically. Due to the stochastic nature of the gradient based optimization process in deep learning with different subsample and mini-batch distributions for each iteration, it is appropriate to use a Probabilistic Graphical Model (PGM) to describe such a stochastic process. PGM has been used to describe sequential decision processes like reinforcement learning (Levine, 2018; Sun and Bischl, 2019). We propose Figure 2 to describe the sequential decision process of model parameter and multiplier adaptation in the form of a probabilistic graphical model (PGM) (Koller and Friedman, 2009). In the upper part of Figure 2, we model the generative process of $N$ observations of the instance $X$ (e.g. an image) and the corresponding supervision signal $Y$ (e.g. a class label). The path $\beta\rightarrow D_{ob}\rightarrow X$, where, $D_{ob}$ is the data-site Sun et al. (2019a) specific prior parameter (domain variable, characterizing distribution) with hyper-prior $\beta$, models the information which is not captured by the supervision signal $Y$ of $X$.
@@
In the lower part of the graphical model, with superscript $k$ representing the iteration for the decision process, we use $\mu^{(k)}$ to represent the multiplier of the loss in Equation 1 at the $k$th iteration. The multiplier $\mu^{(k)}$ serves as a configuration parameter for the optimization process of Equation 1 that affects the value of $\theta_{\mu}^{(k)}$, evolved from its previous value $\theta^{(k-1)}$. $\theta_{\mu}^{(k)}$ then becomes the initial value for the next iteration. This can be described in Equation 14 with explicit dependence of $\theta^{(k)}$ on $\mu^{(k)}$ omitted. The model parameter $\theta^{(k)}$ and observed data $X$, supervision signal $Y$ co-parent the performance indicator as output ${O}^{(k)}$.
@@
An example of $f_{\mu}$ can be based on the acquisition function in Bayesian optimization (BO) (Garnett, 2023) for approximating Equation 3, which leads to a BO realization of this PGM in (Sun et al., 2019a), where $O$ in Equation 11 is usually chosen to be the validation set prediction performance and iteration index $k$ corresponds to a whole training cycle of Equation 2 to iteratively achieve Equation 3.
Different from the BO realization of this PGM, which generates multipliers at a much slower timescale (usually once at the end of training) than that for the evolution of model parameters in neural network training, the hierarchical output feedback optimization we proposed (Figure 1 and elaborated in following sections) operates at the timescale of epoch, thus forms another realization of this PGM with iteration index $k$ corresponding to epochs. In addition, we have $O(\cdot)=\ell(\cdot),R(\cdot)$ for the PGM in Figure 2 in contrast to validation set performance used in BO.
We introduce an objective function for estimating $\theta_{\mu^{(B)}}$ based on the profile likelihood in eq. 12 in accordance with the PGM in Figure 2. This likelihood promotes a high $e\mathcal{HV}(\cdot)$ in Definition 3.1 with respect to $\theta_{\mu^{(B)}}$. $\Xi$ is the normalization factor.
@@
The PGM we defined here can also be used to describe other multi-objective deep learning methods Mahapatra and Rajan (2021) via online combinatorial choice of multipliers.
@@
The PGM also provides a potential possibility to impose a Bayesian interpretation to the joint multiplier tuning and model training process for frequentist machine learning models, the deeper investigation of which we leave for future work.
Figure 3: Illustration to $e\mathcal{HV}$ in Definition 3.1: We take $[\ell(\theta^{(0)},\cdot)\in\mathbb{R},R(\theta^{(0)})\in\mathbb{R}^{d=1}]$ as reference point. Suppose $\{A,B,C,D,E,F\}$ in the illustration constitute $v\mathcal{R}\left(\theta^{(0)},f_{\theta}(\cdot)\right)$ in Definition 2.3. For any $\theta$ corresponding to the point A in the illustration with coordinate $[\ell(\theta,\cdot),R(\theta,\cdot)]$, $e\mathcal{HV}$ maps $\theta$ to $\mathcal{C}_{\ell(\cdot),R(\cdot)}\left(\theta,f_{\theta}(\cdot),\theta^{(0)}\right)$, then to $\mathcal{E}_{\ell(\cdot),R(\cdot)}\left(\theta,f_{\theta}(\cdot),\theta^{(0)}\right)$ (see Definition 2.4), which is the point set {A,B,C,D} , then it calculates the dominated hypervolume with respect to the reference point (union of the shaded rectangles).
3.2 Hierarchical control formulation of joint parameter multiplier optimization

### v3

@@ Abstract
A probabilistic graphical model is proposed, modeling the joint model parameter and multiplier evolution, with a hypervolume based likelihood, promoting multi-objective descent in structural risk minimization. We address multi-objective model parameter optimization via a surrogate single objective penalty loss with time-varying multipliers, equivalent to online scheduling of loss landscape. The multi-objective descent goal is dispatched hierarchically into a series of constraint optimization sub-problems with shrinking bounds according to Pareto dominance. The bound serves as setpoint for the low-level multiplier controller to schedule loss landscapes via output feedback of each loss term. Our method forms closed loop of model parameter dynamic, circumvents excessive memory requirements and extra computational burden of existing multi-objective deep learning methods, and is robust against controller hyperparameter variation, demonstrated on domain generalization tasks with multi-dimensional regularization losses.
@@ 1   Introduction
This work initializes efforts to utilize feedback mechanism, constraint optimization, hierarchical and optimal control into deep learning by treating neural network training as a dynamic process. We tackle two fold challenges in multi-objective deep learning and automatic multiple-multiplier adaptation in structural risk minimization: In many deep learning fields including domain generalization (Gulrajani and Lopez-Paz, 2020), the loss function for neural networks amounts to a summation of multiple terms, with a multiplier weighting each term. As the number of loss terms increases, the space of choice for those multipliers grows dramatically. In addition, when the validation set is drawn from a different distribution than the target domain, the validation performance does not faithfully reflect the target domain performance (Chen et al., 2022b) and can easily reach and maintain a saturated value during training. All these factors complicate and add difficulties to the process of model selection using conventional hyperparameter tuning (Probst et al., 2019).
@@
Inspired by feedback control theory (Doyle et al., 2013) and feedback optimization in power systems (Picallo et al., 2020; He et al., 2023; Hauswirth et al., 2024), we treat the dynamic system of model parameters driven by the low level optimization algorithm as an uncontrolled plant, with the model parameters as state and the many loss terms as output. We use state-dependent multiplier induced loss landscapes (see Figure 4) as control input and designed a hierarchical control structure (Dietterich et al., 1998; Kulkarni et al., 2016; Sun et al., 2020) to drive the system towards a multi-objective descent goal, without modifying the internal dynamic of low level neural network optimization algorithms.
@@
We demonstrate that our method removes the combination curse of dimensionality of multi-dimensional multipliers. Our method only has a small number of hyperparameters for our controllers while achieving robust performance when changing them. We attribute the improved performance of our method to the automatic trade-off of different loss terms.
Our major contributions are:
  • We propose a novel multi-objective optimization algorithm via constraint optimization with shrinking bound. See Proposition 3.1.
  • We propose a probabilistic graphical model for depicting the joint model parameter and multiplier adaptation process with a hypervolume based likelihood, promoting multi-objective descent of each loss term, as an alternative model selection criteria in absence of trustworthy validation data.
  • We bring hierarchical control to solve the joint model parameter and multiplier inference of the probabilistic graphical model as a sequential decision process through an optimal control formulation and optimize the multi-objective descent goal hierarchically by breaking the goal into a series of constraint optimization sub-goals.
@@
  • We bring feedback control theory into deep learning via treating the neural network model parameter dynamic system driven by low level optimization algorithm as an uncontrolled plant. We propose to use state-dependent multiplier-induced loss landscapes as control input and offer initial theoretical analysis of the multi-objective descent behavior of our closed loop system.
@@ 2.1 Structural risk minimization
In Equation 1, we consider loss $L$ with training data $\mathcal{D}_{tr}$ and model parameters $\theta$. $\ell(\theta,\mathcal{D}_{tr})$ represents the empirical risk and $R(\theta,\mathcal{D}_{tr})$ represents the regularization term of the loss function, also referred to as the penalty function, and thus $L$ the penalized loss. $\mathbb{R}_{+}^{d}$ is the $d$ dimensional positive octant. Here $\mu$ is the penalty (weight) multiplier, a special case of general hyperparameters (e.g. learning rate) in deep learning.
2.2 Notation
In the following sections, for brevity, we use $\ell(\cdot)$ and $R(\cdot)$ to indicate the implicit dependence of $\ell$ and $R$ on $\mathcal{D}_{tr}$ and $\theta$. We also write $R(\theta,\cdot)$ where we need an explicit dependence on $\theta$ and use $\cdot$ to implicitly represent other arguments like $\mathcal{D}_{tr}$. We use subscript to indicate the component of $R(\cdot)$, e.g., if $\mu=[\beta,\gamma]$, then $R_{\beta}(\cdot)$ corresponds to the component of $R(\cdot)$ weighted by $\beta$. We use superscript $k$ in bracket to index the optimization iteration (See Remarks 3.1 and 3.2).
Definition 2.1 (Model parameter dynamic system).
When optimizing Equation 1 with multiplier $\mu$ iteratively, we use $\theta^{(k+1)}=f_{\theta}(\mu,\theta^{(k)},\mathcal{D}_{tr},\ell(\cdot),R(\cdot))$ to represent the map bringing $\theta^{(k)}$ to its next value $\theta^{(k+1)}$.
Following $R(\cdot),\ell(\cdot)$, we have $f_{\theta}(\theta,\cdot)$, $f_{\theta}(\mu,\cdot)$ and $f_{\theta}(\cdot)$ to implicitly represent omitted arguments of $f_{\theta}$. We use $\theta^{+}$ for new value of $\theta$ after an operation without explicitly stating how many iterations are needed, resulting in notation $\theta^{+}=f_{\theta}(\mu,\theta,\cdot)$ and $\theta^{+}=f_{\theta}(\theta,\cdot)$.
@@ 2.3 Multi-objective optimization
Definition 2.4 (Non-dominant set map).
If $\theta_{1}\sim\equiv\theta_{2}$, where $\theta_{1}\in s\mathcal{R}(\theta^{(0)},f_{\theta}(\cdot))$ and $\theta_{2}\in s\mathcal{R}(\theta^{(0)},f_{\theta}(\cdot))$ as in Definition 2.3, we define $\theta_{2}\in\mathcal{C}_{\ell(\cdot),R(\cdot)}\left(\theta_{1},f_{\theta}(\cdot),\theta^{(0)}\right)$ the map from a representing element $\theta_{1}$ to its equivalence class, which contains $\theta_{2}$. We use $\mathcal{E}_{\ell(\cdot),R(\cdot)}\left(\theta_{1},f_{\theta}(\cdot),\theta^{(0)}\right)$ to represent the set of $\ell(\cdot),R(\cdot)$ values for each element of $\mathcal{C}_{\ell(\cdot),R(\cdot)}\left(\theta_{1},f_{\theta}(\cdot),\theta^{(0)}\right)$
2.4 Conventional multiplier tuning
The multiplier $\mu$ in Equation 1 is a hyperparameter that affects the evolution of the model parameters $\theta$ in the optimization process for $L$ defined in Equation 1, which leads to optimized model parameter $\theta_{\mu,\mathcal{D}_{tr}}$ in Equation 2, where we omit the potential dependence of $\theta_{\mu,\mathcal{D}_{tr}}$ on its initial condition $\theta^{(0)}$ for notational simplicity.
@@ 2.4 Conventional multiplier tuning
Hyperparameter optimization aims to adjust the multiplier $\mu$ in alignment with specific performance metrics ${O}$ (e.g. validation set accuracy for a classification task) in Equation 3 evaluated on the validation set $\mathcal{D}_{val}$, which leads to the selected hyperparameter $\mu_{\mathcal{D}_{tr},\mathcal{D}_{val}}$.
@@
To solve Equation 3 iteratively, each iteration relies on a complete training cycle to optimize Equation 2. The selection of $\mu$ depends on validation set $\mathcal{D}_{val}$, which can only be drawn from a different distribution compared to the target domain for the domain generalization problem (Section C.2). Additionally, the performance metric $O$ on the in-domain validation set can reach and maintain a saturated value during training with neural networks of high expressive power.
@@ 3   Methods
In this section, we elaborate on our multi-objective hierarchical output feedback optimization (M-HOF-Opt) as a control strategy in Figure 1 for the multidimensional multiplier adaptation in Equation 1. We propose a probabilistic graphical model for the joint model parameter and multiplier adaptation process in Section 3.1, which leads to an optimal control formulation of parameter estimation, extending the classical hyperparameter tuning-based approach to parameter-multiplier co-evolution in Section 3.2.
@@
Figure 1: Control diagram illustrating the hierarchical output feedback optimization process with multi-objective setpoint adaptation. The uncontrolled plant corresponds to an open loop dynamic $\theta^{+}=f_{\theta}(\mu,\theta,\cdot)$ defined in Definition 2.1, via optimizing Equation 1 with a low level optimization algorithm lacking feedback. The $\mu$ controller adjusts the multiplier $\mu$, thus schedules the loss landscape as illustrated in Figure 4, based on the difference $e$ between the setpoint $b$ and the measured output components $R(\cdot)$, guiding the optimization of model parameters $\theta$ through a feedback loop. The setpoint $b$ is adjusted via feedback from $\ell(\cdot)$ and $R(\cdot)$, which forms a higher hierarchy. See Figure 2 for the probabilistic description of the closed loop behavior of this control diagram.
3.1 Probabilistic graphical model for joint decision of model parameter and multiplier
@@
Gradient based optimization algorithms for neural networks define a dynamic system for model parameters where the weight multiplier in Equation 1 configures such a dynamic system via defining a loss landscape (see Figure 4). Due to the stochastic nature of the gradient with different subsample and mini-batch distributions for each iteration, it is appropriate to use a probabilistic graphical model (Koller and Friedman, 2009) to describe such a stochastic process, which has been used to describe sequential decision processes like reinforcement learning (Levine, 2018; Sun and Bischl, 2019). Instead of fixed multipliers, we also model multiplier choice adaptively and propose the probabilistic graphical model in Figure 2 to describe the sequential decision process of the joint model parameter and multiplier adaptation.
Figure 2: Probabilistic graphical model for the sequential decision process of joint model parameter and multiplier adaptation in multi-domain structral risk minimization in Equation 1. See Figure 1 for the control diagram counterpart of the same process.
In the upper part of Figure 2, we model the generative process of $N$ observations of $X$ (e.g. an image) and the corresponding supervision signal $Y$ (e.g. a class label or self-supervision). The path $\beta\rightarrow D_{ob}\rightarrow X$ models domain-specific data generation (information not captured by $Y$), where $D_{ob}$ modeling the data-site Sun et al. (2019a) with hyper-prior $\beta$.
@@
In the lower part of Figure 2, with superscript $k$ indexing the iteration of the decision process, we use $\mu^{(k)}$ to represent the multiplier of the loss in Equation 1 at the $k$th iteration. The multiplier $\mu^{(k)}$ serves as a configuration parameter for the optimization process of Equation 1 that affects the value of $\theta_{\mu}^{(k)}$ (evolved from its previous value $\theta^{(k-1)}$). $\theta_{\mu}^{(k)}$ then becomes the initial value for the next iteration. This can be described in Equation 7. The model parameter $\theta^{(k)}$ and observed data $X$, supervision signal $Y$ co-parent the performance indicator ${O}^{(k)}$ as output of the dynamic system. See different realizations of $O$ in Remark 3.1 and Remark 3.2.
@@
A realization of $f_{\mu}$ can be based on the acquisition function in Bayesian optimization (Garnett, 2023) for approximating Equation 3. See an example described in Sun et al. (2019a), where $O$ in Equation 4 is chosen to be the validation set prediction performance and iteration index $k$ corresponds to a whole training cycle of Equation 2 to iteratively achieve Equation 3.
Remark 3.2.
Different from the Bayesian optimization realization of this probabilistic graphical model, which changes multipliers at a much slower timescale than that for the evolution of model parameters in neural network training, the hierarchical output feedback optimization we proposed in Figure 1 and elaborated in Section 3.2 and following sections operates at the timescale of epoch, thus forms another realization of this probabilistic graphical model with the iteration index $k$ corresponding to an epoch. In addition, we have $O(\cdot)=\ell(\cdot),R(\cdot)$ in Figure 2 in contrast to validation set performance used in Bayesian optimization.
If the joint decision process of model parameter and multiplier is optimal, from a statistical inference point of view (See Remark 3.5), we are conducting a sequential inference (i.e., finding a sequence of $\mu$ in Equation 6), such that at the last step $B$, the final value $\theta_{\mu^{(B)}}$ has the best possible hyper-volume (in a stochastic sense, see Figure 3) with respect to output $O$ (see Remark 3.1 and Remark 3.2).
We introduce an objective function for estimating $\theta_{\mu^{(B)}}$ based on the profile likelihood in eq. 5 in accordance with the probabilistic graphical model in Figure 2. This likelihood promotes a high $e\mathcal{HV}(\cdot)$ in Definition 3.1 with respect to $\theta_{\mu^{(B)}}$. $\Xi$ is the normalization factor.
@@
Figure 3: Illustration of $e\mathcal{HV}$ in Definition 3.1: We take $[\ell(\theta^{(0)},\cdot),R(\theta^{(0)})]$ as reference point. Suppose $\{A,B,C,D,E,F\}$ in the illustration constitutes function values of the reachable set $v\mathcal{R}\left(\theta^{(0)},f_{\theta}(\cdot)\right)$in Definition 2.3. For any $\theta$ corresponding to the point A in the illustration with coordinate $[\ell(\theta,\cdot),R(\theta,\cdot)]$, $e\mathcal{HV}$ maps $\theta$ to $\mathcal{C}_{\ell(\cdot),R(\cdot)}\left(\theta,f_{\theta}(\cdot),\theta^{(0)}\right)$, then to $\mathcal{E}_{\ell(\cdot),R(\cdot)}\left(\theta,f_{\theta}(\cdot),\theta^{(0)}\right)$ (see Definition 2.4), which is the point set {A,B,C,D}, then it calculates the dominated hypervolume with respect to the reference point (union of the shaded rectangles).
@@
The probabilistic graphical model in Figure 2 can also be used to describe other multi-objective deep learning methods (Mahapatra and Rajan, 2021) via online combinatorial choice of multipliers.
Remark 3.4.
The probabilistic graphical model also provides a potential possibility to impose a Bayesian interpretation to the joint model parameter training process and multiplier tuning for frequentist machine learning, the deeper investigation of which we leave for future work.
Remark 3.5 (The dual relationship between estimation and control).
Analogous to how optimal control Lewis et al. (2012) can be formulated as maximum likelihood estimation in a probabilistic graphical model (Levine, 2018; Sun and Bischl, 2019), the estimation of $\theta^{(B)}$ based on the profile likelihood like objective in Equation 5 for Figure 2 is also a sequential decision process and can be solved via optimal control.
3.2 Hierarchical control formulation of joint model parameter and multiplier optimization